# HAFO: Humanoid Force-Adaptive Control for Intense External Force Interaction Environments

### v2

@@ Abstract
Reinforcement learning (RL) controllers have made impressive progress in humanoid locomotion and light load manipulation. However, achieving robust and precise motion with strong force interaction remains a significant challenge. Based on the above limitations, this paper proposes HAFO, a dual-agent reinforcement learning control framework that simultaneously optimizes both a robust locomotion strategy and a precise upper-body manipulation strategy through coupled training under external force interaction environments. Simultaneously, we explicitly model the external pulling disturbances through a spring-damper system and achieve fine-grained force control by manipulating the virtual spring. During this process, the reinforcement-learning policy spontaneously generates disturbance-rejection response by exploiting environmental feedback. Moreover, HAFO employs an asymmetric Actor-Critic framework in which the Critic-network access to privileged spring-damping forces guides the actor-network to learn a generalizable, robust policy for resisting external disturbances. The experimental results demonstrate that HAFO achieves stable control of humanoid robot under various strong force interactions, showing remarkable performance in load tasks while ensuring stable robot operation under rope tension disturbances.
@@ 1 Introduction
In recent years, RL-based control solutions have made significant progress in the field of humanoid robot whole-body control [1, 2, 3, 4, 5, 6]. Leveraging large-scale parallel simulations and domain randomization, policies are trained to absorb disturbances in mass, friction, and contact parameters, enabling zero-shot transfer to the real robot. However, RL controllers often neglect or oversimplify interactive forces. To meet the practical needs of humanoid robots in tasks like heavy-load manipulation and high-altitude operation, precise and robust motion control is essential in environments with strong external forces and dynamic changes.
When humanoid robots subject a strong collision or physical contact, RL controllers often exhibit instability. Specifically, current RL methods primarily use high-level commands such as moving velocity, root height, or predefined trajectories for control, which can overcome some random disturbances. However, when subjected to human intervention or significant contact with the environment, the lack of explicit modeling of contact dynamics often makes it difficult for the robot to maintain balance, potentially causing safety hazards [7, 8]. For example, in high-altitude operations, ropes are required for assistance, but their interference significantly affects the stability of the robot. In heavy-load scenarios, large payloads greatly reduce balanced behavior, placing severe demands on the load-bearing capacity of the hand.
To enhance the stability of control strategies under human intervention or significant external disturbances, this paper introduces HAFO, a force-adaptive control framework grounded in explicit dynamic modeling. HAFO employs a dual-agent strategy that decouples the control of the upper and lower bodies. The lower-body strategy focuses on robust locomotion to maintain gait stability under strong external disturbances. The upper-body strategy emphasizes precise control and real-time upper-body manipulation under external force disturbances. We utilize a spring-damping mechanical system to simulate interaction mechanics modeling. Meanwhile, we progressively increase external forces through a curriculum-learning schedule and randomize the point of force application, systematically enhancing the strategy’s generalization to pulling forces, heavy load, and thrust disturbances. Our contributions are as follows:
@@
    A dual-agent reinforcement learning framework enabling the collaborative evolution of the upper and lower bodies under large external disturbances is proposed. The lower-body strategy is responsible for robust locomotion, and the upper-body strategy focuses on precise manipulation. Both strategies share the same state space. Through a coupling-adversarial mechanism, the system’s stability is significantly improved, ultimately achieving whole-body collaborative control.
@@
    A Spring-Damping dynamic model has been developed, which explicitly models the external pulling force, as equivalent spring-damper forces acting on the robot body. During model training, a carefully designed progressive curriculum learning strategy is implemented, gradually increasing the spring-damper forces, enabling the policy gradually adapt to external disturbances. The RL policy autonomously evolves force-adaptive control modes for ground locomotion and aerial suspension, and spontaneously generates mode-switching behaviors based on environmental feedback, without relying on explicit state machines or predefined switching logic.
@@
    Experimental results demonstrate that the HAFO policy enables adaptive force control across multiple environments, exhibiting stable and precise motion capabilities under single or dual-hand loads. Moreover, HAFO is the first locomotion-control strategy to achieve stable operation and safe startup under rope-suspension conditions, representing preliminary exploratory work for humanoid robot applications in high-altitude scenarios.
@@ 2.1 Whole-body control based on reinforcement learning
Object handling and assembly tasks demand whole-body coordination of a humanoid, which drastically increases control complexity. Traditional dynamics modeling has demonstrated good control performance in structured scenarios for legged robots [9, 10, 11, 12, 13, 14, 15, 16, 17]. However, the intrinsic under-actuation of humanoid robots and the variability of real-world environments make it challenging to accurately describe the dynamics of humanoid robots [1, 2, 3, 4, 6, 7, 8, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32]. Recently, some technical solutions based on reinforcement learning provide new ideas for the whole-body control of humanoid robots. Common approaches include Lower-RL-Upper-IK and Monolithic-Whole-body-RL. These two paradigms solve the challenge of whole-body motion control of humanoid robots from different perspectives.
@@
In terms of Lower-RL-Upper-IK paradigm, Ben et al. [18] proposed the Homie, which employs curriculum learning to dynamically adjust the motion tracking threshold, allowing the policy to progressively adapt to high-difficulty motions during training. Additionally, the Homie controller develops an isomorphic exoskeleton cockpit to achieve precise whole-body control of the robot. Liao et al. [27] designed a scalable and high-quality motion tracking framework called BeyondMimic, which can transform kinematic references into robust and highly dynamic robot motions. Through distillation, it integrates the trained motions into a single policy, enabling flexible combination of multiple skills during testing and achieving goal-driven control using only simple cost functions. By combining large-scale human motion data with reinforcement learning, Cheng et al. [19] encourage the robot’s upper body to imitate the reference motion, while relaxing the imitation constraints on its two legs, enabling the robot to achieve expressive movements. The Mobile-TeleVision proposed by Lu et al. [23] introduces the predictive motion prior (PMP), which encodes multi-modal motion features through conditional variational autoencoder (CVAE) to guide the reinforcement learning strategy generates dynamic motions. For the Monolithic-Whole-body-RL paradigm, Xie et al. [28] proposed a KungfuBot, which introduces foot-ground contact states and a motion-phase signal to accurately imitate highly dynamic wushu movements. He et al. [2] proposed the OmniH2O teleoperation system, which integrates a teacher-student distillation framework to transfer knowledge from a privileged teacher policy to the deployed policy, enabling efficient and stable motion control of the robot in real-world environments. He et al. [21] proposed ASAP, which significantly enhances a robot’s adaptability in complex environments by leveraging a pretrained motion tracking policy with a residual action model trained using real-world data.
@@
Although the humanoid robot whole-body motion control has made remarkable progress in the above research, the existing methods still face many challenges. On the one hand, the Lower-RL-Upper-IK paradigm adopts the control mode of the upper-body open loop, which is difficult to perceive and adapt to the external interference in real time, on the other hand, the Monolithic-Whole-body-RL paradigm controls whole-body joints through a single controller, but due to the weak task goal correlation between upper-limb manipulation and lower-limb walking, it is prone to overfitting.
@@ 2.2 Force-adaptive control
Humanoid robots inevitably experience continuous disturbances from external forces in real-world environments. However, existing research on force-adaptive control is relatively scarce. In the field of manipulation, hybrid position and force control methods are widely used. However, given the uncertainty of the interaction points between humanoid robots and the environment, the interaction forces are difficult to measure accurately using force sensors, which makes it challenging to directly apply these methods to legged robots.
Portela [33] et al. first introduced the force-adaptive control objective for quadruped robots, which significantly improved the force interaction capability of the end effector. However, the model requires explicit switching between position mode and force mode, and assumes the system is in a quasi-static state, making it difficult to handle dynamic and unknown disturbances. Force-adaptive control in humanoid robots presents several challenges. Specifically, there is significant dynamic coupling between the upper and lower limbs, joint torques are constrained by strict limitations, and the mass distribution is uneven. These characteristics make force-adaptive control particularly complex [34, 35, 36]. It is particularly noteworthy that FALCON designed a 3D force curriculum, which maximizes the force adaptability while ensuring the feasibility of joint torques [6]. Li et al. [37] introduced a force-adaptive torso-tilt reward, which maintains balance and effectively transmits force through torso tilting and center-of-gravity adjustment, generating greater pulling forces in high-intensity force interaction tasks. Kuo et al. [38] proposed a hybrid control method combining the Artificial Rabbit Optimization (ARO) algorithm with an MLP neural network. Through multi-angle impact experiments on flat and sloped terrain, they verified that the method can quickly adapt to external forces and effectively maintain balance. Existing model-based methods rely on precise modeling of pre-planned force trajectories, which restricts their generalization capability in unknown environments, rendering them inadequate for adapting to dynamically changing conditions [39]. In contrast, estimation-based approaches [40] are only applicable under quasi-static conditions and fail to effectively address force adaptive control in dynamic loco-manipulation tasks.
Although existing studies have explored force adaptive control for humanoid robots, most focus on relatively simple scenarios. Moreover, studies implementing force adaptive control on humanoid robots are notably sparse. In this paper, we propose a force adaptation whole-body control (HAFO) strategy. Through explicit modeling of external disturbances, the proposed strategy endows humanoid robots with force-adaptive capability under challenging conditions such as heavy loads and pulling disturbances.
@@ 3 Method
Humanoid robot motion control requires coordinated movement of whole-body joints. Traditional whole-body control methods (WBC) use a single strategy to output the target joint angles for the entire body. However, the lack of strong coupling between upper and lower bodies tasks leads to inefficient sample utilization, making prone to overfitting. To maximize the exploration efficiency of reinforcement learning, this paper proposes the HAFO architecture shown in Fig. 1, in which a dual-agent reinforcement-learning framework drives the co-evolution of upper and lower bodies control strategies. Moreover, during the training process, a variety of dynamic disturbance factors are explicitly introduced to achieve stable force-adaptive control.
Figure 1: Overview of the HAFO model. (a) Policy training. A dual-agent strategy with decoupled upper and lower bodies is adopted, where the lower-body policy takes root linear and angular velocities as command inputs, and the upper-body policy uses reference joint trajectories as command inputs. Meanwhile, various explicit dynamic perturbations are introduced at key locations to enhance the system’s robustness and adaptability. (b) Strategy deployment. A humanoid robot control system based on teleoperation is developed, employing an efficient inverse kinematics algorithm to compute the robot’s joint angles in real time with high precision, enabling efficient loco-manipulation tasks.
@@ 3.1 Dual-agent learning framework
We decompose the total DOFs of the humanoid into upper and lower bodies, $n_{\mathrm{dof}}=\mathit{\text{Upper}}_{\mathrm{\text{dof}}}+\mathit{\text{Lower}}_{\mathrm{dof}}$, where $\mathit{\text{Upper}}_{\text{dof}}$ is controlled by upper-body strategy $\pi_{\text{Upper}}\left(\mathrm{a}_{\text{Upper }}\mid\mathrm{s}\right)$, $\mathit{\text{Lower}}_{\text{dof}}$ is controlled by the lower-body strategy $\pi_{\text{lower }}\left(\mathrm{a}_{\text{lower }}\mid\mathrm{s}\right)$. The dual-agent reinforcement learning strategy is trained using the Proximal Policy Optimization (PPO) algorithm [41] and shares the same state space ${s}_{{t}}=\left[{s}_{{t}}^{\text{prop }},\mathrm{c}^{l},\mathrm{~g}^{{u}}\right]$, enabling both agents to consider the information of whole-body joints and achieve coordinated control. Among them, the ${s}_{\mathrm{t}}^{\text{prop }}=\left[{q}_{{t}},\dot{q}_{t},\omega_{t},gv_{t},a_{t-1}\right]$ represents the body perception information of the robot, the $c^{l}=\left[\hat{v}_{x,t},\hat{v}_{y,t},\hat{\omega}_{yaw},t\right]$ represents the command information for the lower-body strategy, and the ${g}^{{u}}\in\mathbb{R}^{14}$ represents the command information for the upper-body strategy.
The dual-agent policy is formulated as a dual-objective optimization problem. The value functions $\left(V^{l},V^{u}\right)$ corresponding to the upper and lower body policies $\left(\pi^{l},\pi^{u}\right)$, and the rewards obtained can be represented as $\left(r^{l}\left(s,a^{l},a^{u}\right),r^{u}\left(\mathrm{~s},\mathrm{a}^{1},\mathrm{a}^{\mathrm{u}}\right)\right)$. The two strategies optimize the policies by maximizing the cumulative rewards [42],
@@ 3.2 Robust lower-body movement under external forces
Robust locomotion of humanoid robots is still a formidable challenge, primarily due to the complexity of the floating base and the instability caused by high inertia. This difficulty is compounded when payloads are placed on the upper body, which both increases the overall inertia and introduces additional modeling uncertainty, further complicating control design.
For the lower limb strategy $\pi^{l}:C^{l}\times S\mapsto A^{l}$, we employ a variety of external disturbances during the training process to enhance the robot’s robustness, including adversarial upper-body perturbations and high-frequency random pushing disturbances.

### v3

# HAFO: A Force-Adaptive Control Framework for Humanoid Robots in Intense Interaction Environments

@@ Abstract
Reinforcement learning (RL) controllers have made impressive progress in humanoid locomotion and light-weight object manipulation. However, achieving robust and precise motion control with intense force interaction remains a significant challenge. To address these limitations, this paper proposes HAFO, a dual-agent reinforcement learning framework that concurrently optimizes both a robust locomotion strategy and a precise upper-body manipulation strategy via coupled training in environments with external disturbances. The external pulling disturbances are explicitly modeled using a spring-damper system, allowing for fine-grained force control through manipulation of the virtual spring. In this process, the reinforcement learning policy autonomously generates a disturbance-rejection response by utilizing environmental feedback. Furthermore, HAFO employs an asymmetric Actor-Critic framework in which the Critic network’s access to privileged external forces guides the actor network to acquire generalizable force adaptation for resisting external disturbances. The experimental results demonstrate that HAFO achieves whole-body control for humanoid robots across diverse force-interaction environments, delivering outstanding performance in load-bearing tasks and maintaining stable operation even under rope suspension state.
@@ 1 Introduction
In recent years, learning-based control solutions have made significant progress in the field of whole-body control for humanoid robots [1, 2, 3, 4, 5, 6]. By leveraging large-scale parallel simulations and domain randomization, policy networks can effectively learn to handle environmental uncertainties, thus enabling zero-shot transfer to real robots. However, RL controllers often overlook or oversimplify interactive forces. To address the practical needs of humanoid robots in tasks such as heavy-load manipulation and high-altitude operations, precise and robust motion control is crucial in environments with strong external forces and dynamic changes.
When humanoid robots encounter strong collisions or physical contact, reinforcement learning (RL) controllers often exhibit instability. Specifically, current RL methods primarily rely on high-level commands, such as movement velocity, root height, or predefined trajectories, which may mitigate some random disturbances. However, when exposed to human intervention or significant environmental contact, the absence of explicit modeling of contact dynamics often hinders the robot’s ability to maintain balance, potentially leading to safety hazards [7, 8]. For example, in high-altitude operations, ropes are required for support, but their interference can significantly undermine the robot’s stability. In heavy-load scenarios, large payloads significantly impair balance, placing considerable demands on the load-bearing capacity.
To enhance the stability of control strategies under human intervention or significant external disturbances, this paper introduces HAFO, a force-adaptive control framework based on explicit dynamic modeling. HAFO employs a dual-agent strategy that decouples the control of the upper and lower bodies. The lower-body strategy focuses on robust locomotion to maintain gait stability under strong external disturbances. The upper-body strategy emphasizes precise and real-time upper-body manipulation under external force disturbances. We utilize a spring-damper mechanical system to model external pulling disturbances. Meanwhile, we progressively increase external forces through a curriculum-learning schedule and randomize the point of force application, systematically enhancing the strategy’s generalization to various pulling disturbances and different hand loads. The contributions of this paper are as follows:
@@
    A dual-agent reinforcement learning framework is proposed to facilitate the collaborative evolution of the upper and lower bodies in the presence of large external disturbances. The lower-body strategy focuses on ensuring robust locomotion, and the upper-body strategy focuses on precise manipulation. Both strategies share a common state space. Through a coupling-adversarial mechanism, the stability of the entire body system under disturbances is significantly improved, enhancing the robot’s control robustness against multi-source external disturbances.
@@
    A spring-damping dynamic model is developed to explicitly model the external pulling force as equivalent spring-damper forces acting on the robot’s body. During model training, a carefully designed progressive curriculum learning strategy is employed, gradually increasing the spring-damper forces to enable the policy to adapt progressively to external disturbances. The reinforcement learning (RL) policy autonomously evolves force-adaptive control modes for ground locomotion and aerial suspension, spontaneously generating mode-switching behaviors based on environmental feedback, without relying on explicit state machines or predefined switching logic.
@@
    Experimental results demonstrate that the HAFO policy enables adaptive force control across various environments, exhibiting stable and precise motion capabilities under single or dual-hand loads. Moreover, HAFO is the first locomotion-control strategy to achieve stable operation and safe startup under rope-suspension conditions, marking preliminary exploratory work for humanoid robot applications in high-altitude environments.
@@ 2.1 Whole-body control based on reinforcement learning
Conventional dynamic modeling methods have shown impressive control performance in structured scenarios for legged robots [9, 10, 11, 12, 13, 14, 15, 16, 17]. However, due to the intrinsic under-actuation of humanoid robots and the variability of real-world environments, it is extremely difficult to establish an accurate dynamics model. Recently, several RL-based solutions have provided new insights for whole-body control of humanoid robots [1, 2, 3, 4, 6, 7, 8, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32]. Common paradigms include Lower-RL-Upper-IK and Integrated Whole-body RL, which address the challenge of whole-body motion control from different perspectives.
@@
In terms of Lower-RL-Upper-IK paradigm, Ben et al. [18] proposed the Homie, which employs curriculum learning to dynamically adjust the motion tracking threshold, allowing the policy to progressively adapt to high-difficulty motions during training. Additionally, the Homie controller develops an isomorphic exoskeleton cockpit to achieve precise whole-body control of the robot. By combining large-scale human motion data with reinforcement learning, Cheng et al. [19] encourage the robot’s upper body to imitate the reference motion, while relaxing the imitation constraints on the legs, allowing the robot to achieve expressive movements. The Mobile-TeleVision proposed by Lu et al. [23] introduces the predictive motion prior (PMP), which encodes multi-modal motion features through conditional variational autoencoder (CVAE) to guide the reinforcement learning strategy generates dynamic motions. For the Integrated Whole-body RL paradigm, Xie et al. [28] proposed a KungfuBot, which incorporates foot-ground contact states and a motion-phase signal to accurately imitate highly dynamic wushu movements. He et al. [2] introduced the OmniH2O teleoperation system, which integrates a teacher-student distillation framework to transfer knowledge from a privileged teacher policy to the deployed policy, enabling efficient and stable motion control of the robot in real-world environments. He et al. [21] proposed ASAP, which significantly enhances a robot’s adaptability in complex environments by leveraging a pretrained motion tracking policy with a residual action model trained using real-world data. Liao et al. [27] designed a scalable and high-quality motion tracking framework called BeyondMimic, which transforms kinematic references into robust, highly dynamic robot motions. Through distillation, it integrates the trained motions into a single policy, enabling the flexible combination of multiple skills during testing and achieving goal-driven control using simple cost functions.
@@
Although whole-body motion control for humanoid robots has made remarkable progress in the research mentioned above, the existing methods still face many challenges. On the one hand, the Lower-RL-Upper-IK paradigm employs an open-loop control mode for the upper body, which struggles to perceive and adapt to external interference in real time. On the other hand, the Integrated paradigm controls whole-body joints through a single controller, but due to the weak task goal correlation between upper-body manipulation and lower-body walking, it is prone to overfitting.
@@ 2.2 Force-adaptive control
Humanoid robots inevitably experience continuous disturbances from external forces in real-world environments. However, existing research on force-adaptive control is relatively scarce. In the field of manipulation, hybrid position and force control methods are commonly employed. Nevertheless, due to the uncertainty of interaction points between humanoid robots and the environment, accurately measuring interaction forces with force sensors becomes challenging, making it difficult to directly apply these methods to legged robots.
Portela [33] et al. first introduced the force-adaptive control objective for quadruped robots, which significantly improved the force interaction capability of the end effector. However, the model requires explicit switching between position mode and force mode, and assumes the system is in a quasi-static state, making it difficult to handle dynamic and unknown disturbances. For humanoid robots, their inherent characteristics, such as a high center of gravity and a small support base, make force-adaptive control particularly complex [34, 35, 36]. Li et al. [37] introduced a force-adaptive torso-tilt reward, which maintains balance and effectively transmits force through torso tilting and center-of-gravity adjustment, generating greater pulling forces in high-intensity force interaction tasks. Zhang et al. [6] designed a 3D force curriculum, which maximizes the force adaptability while ensuring the feasibility of joint torque, achieving high-intensity force interactions. Kuo et al. [38] proposed a hybrid control method combining the Artificial Rabbit Optimization (ARO) algorithm with an MLP neural network. Through multi-angle impact experiments on flat and sloped terrain, they verified that the method can quickly adapt to external forces and effectively maintain balance. Existing model-based methods rely on precise modeling of pre-planned force trajectories, which restricts their generalization capability in unknown environments, rendering them inadequate for adapting to dynamically changing conditions [39]. In contrast, estimation-based approaches [40] are only applicable under quasi-static conditions and fail to effectively address force adaptive control in dynamic loco-manipulation tasks.
Although existing studies have explored force adaptive control for humanoid robots, the range of scenarios explored so far remains relatively limited. In this paper, we propose a force-adaptive whole-body control strategy (HAFO), which endows humanoid robots with force-adaptive capabilities in challenging conditions, such as heavy loads and pulling disturbances.
@@ 3 Method
Humanoid robot motion control requires coordinated movement of whole-body joints. Traditional whole-body control methods rely on a single strategy to output target joint angles for the entire body. However, the lack of strong coupling between upper and lower bodies tasks leads to inefficient sample utilization, making the policy prone to overfitting. To maximize the exploration efficiency of reinforcement learning, this paper proposes the HAFO architecture shown in Figure. 1, which employs a dual-agent reinforcement learning framework to drive the co-evolution of upper and lower bodies control strategies. Additionally, during training, various dynamic disturbances are explicitly introduced to ensure stable force-adaptive control.
Figure 1: Overview of the HAFO model. (a) Policy Training. A dual-agent strategy with decoupled upper and lower bodies is adopted, where the lower-body policy takes root linear and angular velocities as command inputs, and the upper-body policy uses reference joint trajectories as command inputs. Additionally, various explicit dynamic perturbations are introduced at key locations to enhance the system’s robustness and adaptability. (b) Strategy Deployment. A humanoid robot control system based on teleoperation is developed, employing an efficient inverse kinematics algorithm to compute the robot’s joint angles in real time with high precision, enabling efficient loco-manipulation tasks.
@@ 3.1 Dual-agent learning framework
We decompose the total degrees of freedom (DOFs) of the humanoid into upper and lower bodies, such that $n_{\mathrm{dof}}=\mathit{\text{Upper}}_{\mathrm{\text{dof}}}+\mathit{\text{Lower}}_{\mathrm{dof}}$, where $\mathit{\text{Upper}}_{\text{dof}}$ is controlled by upper-body strategy $\pi_{\text{Upper}}\left(\mathrm{a}_{\text{Upper }}\mid\mathrm{s}\right)$, $\mathit{\text{Lower}}_{\text{dof}}$ is controlled by the lower-body strategy $\pi_{\text{lower }}\left(\mathrm{a}_{\text{lower }}\mid\mathrm{s}\right)$. The dual-agent reinforcement learning strategy is trained using the Proximal Policy Optimization (PPO) algorithm [41] and shares the same state space ${s}_{{t}}=\left[{s}_{{t}}^{\text{prop }},\mathrm{c}^{l},\mathrm{~g}^{{u}}\right]$, allowing both agents to consider information from the entire body and achieve coordinated control. Among them, the ${s}_{\mathrm{t}}^{\text{prop }}=\left[{q}_{{t}},\dot{q}_{t},\omega_{t},gv_{t},a_{t-1}\right]$ represents the body perception information of the robot, the $c^{l}=\left[\hat{v}_{x,t},\hat{v}_{y,t},\hat{\omega}_{yaw},t\right]$ represents the command information for the lower-body strategy, and the ${g}^{{u}}\in\mathbb{R}^{14}$ represents the command information for the upper-body strategy.
The dual-agent policy is formulated as a dual-objective optimization problem. The value functions $\left(V^{l},V^{u}\right)$ corresponding to the upper and lower body policies $\left(\pi^{l},\pi^{u}\right)$, and the rewards obtained can be represented as $\left(r^{l}\left(s,a^{l},a^{u}\right),r^{u}\left(\mathrm{~s},\mathrm{a}^{1},\mathrm{a}^{\mathrm{u}}\right)\right)$. The dual policies are optimized by maximizing the cumulative rewards [42],
@@ 3.2 Robust lower-body movement under external forces
Robust locomotion of humanoid robots remains a significant challenge, primarily due to the complexity of the floating base and the instability caused by high inertia. This difficulty is further compounded when payloads are applied to the upper body, increasing the overall inertia and introducing additional modeling uncertainties, which complicate the control design.
For the lower-body strategy $\pi^{l}:C^{l}\times S\mapsto A^{l}$, various external disturbances are introduced during the training process to enhance the robot’s robustness, including adversarial upper-body perturbations and high-frequency random pushing disturbances.